OpenStudy (anonymous):

72 is what precent of 400? Write and solve an equaion to solve the problem

OpenStudy (anonymous):

18%

OpenStudy (anonymous):

Use the names of operations to convert this to an algebraic equation: "is" means "=" Percent means "/100" (divided by 100) "of" means "*" (multiply) "what number" means to insert your favorite letter for a variable (usually the letter x) 72=18/100 * x 72 = .18x 72/.18 = x 400 = x Mental Check: 10% of 400 is 1/10 of 400 which is 40. 20% of 400 is 2/10 of 400 which is 80. Is 72 larger than 40, less than 80 but closer to 80? Yes. So I think we're at least close. Use your calculator: 18% of 400 or .18 * 400 = 72!!